OpenStudy (anonymous):

Find the sum of the geometric sequence. 4/3, 8/3, 16/3, 32/3, 64/3

OpenStudy (anonymous):

@satellite73 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

What's the first term? What's the common ratio?

OpenStudy (anonymous):

4/3

OpenStudy (anonymous):

1/2?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

4/3 is the correct first term

jimthompson5910 (jim_thompson5910):

1/2 is not the common ratio

jimthompson5910 (jim_thompson5910):

1/2 is close though

OpenStudy (anonymous):

1/3? hahs

jimthompson5910 (jim_thompson5910):

solve for r (4/3)r = 8/3

OpenStudy (anonymous):

2 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Now use the formula \[\Large S_{n} = a*\frac{1-r^n}{1-r}\]

jimthompson5910 (jim_thompson5910):

in this case a = 4/3 r = 2 n = 5

OpenStudy (anonymous):

i got -44

jimthompson5910 (jim_thompson5910):

that's incorrect

jimthompson5910 (jim_thompson5910):

\[\Large S_{n} = a*\frac{1-r^n}{1-r}\] \[\Large S_{5} = \left(\frac{4}{3}\right)*\frac{1-2^5}{1-2}\] \[\Large S_{5} = ???\]

OpenStudy (anonymous):

-32.6666666667

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

that's not correct either

jimthompson5910 (jim_thompson5910):

what is 1 - 2^5 ?

OpenStudy (anonymous):

-31

jimthompson5910 (jim_thompson5910):

1 - 2 = ???

OpenStudy (anonymous):

-1

jimthompson5910 (jim_thompson5910):

\[\Large S_{n} = a*\frac{1-r^n}{1-r}\] \[\Large S_{5} = \left(\frac{4}{3}\right)*\frac{1-2^5}{1-2}\] \[\Large S_{5} = \left(\frac{4}{3}\right)*\frac{-31}{-1}\] \[\Large S_{5} = ???\]

OpenStudy (anonymous):

41.3333333333

jimthompson5910 (jim_thompson5910):

or 124/3

OpenStudy (anonymous):

Thank you♥

jimthompson5910 (jim_thompson5910):

you're welcome